OpenStudy (anonymous):

how many solutions will the following system of linear equations have? -2x+4y=-7 y=-1/2x+5 A.0 B.1 C.2 D.infinite

OpenStudy (anonymous):

-2x+4y=-7 y=-1/2x+5 substitute y=-1/2x+5 into the first equation wherever you see y. -2x + 4*(-1/2x+5) = -7 -2x -2x + 20 = -7 -4x = 13 x = -13/4 plug this value into either equation to find y.

OpenStudy (anonymous):

-2x -2x + 20 = -7 -4x = -27 x = 27/4

OpenStudy (anonymous):

i messed up on the subtraction

OpenStudy (e.mccormick):

Do you see what Tomo was doing? If you repeat it, but without the mistake, you can get an answer. However, there is another way to think about it for that particular question. because this is a linear system, the choices are 0, 1, and \(\infty\) solutions. They are represented by the 3 ways the lines can be in space.  The slope can tell you that.